OpenStudy (anonymous):

Was this done correctly?? -4x+3(1*12)≥40+2*8 -4x+3*36≥40+2*8 -4x+108≥40+2*8 -4x+108≥40+2*8 -4x+108≥40+16 -4x+108≥56 -4x+108-108≥56-108 -4x≥-52 -4x/-4≥-52/-4 x≥13

OpenStudy (anonymous):

Will medal

OpenStudy (loser66):

nope

OpenStudy (loser66):

line1 to line2, the second term of the left hand side: why 3(1*12) becomes 3*36?

OpenStudy (anonymous):

because 3*1 is 3 and 3*12 is 36..distributive property right?

OpenStudy (loser66):

nope, if it is 3(1+12) then it is 3*1 +3*12 , but it is 3(1* 12) and `1*12 =12 only , hence 3(1*12) =3*12 = 36, not 3*36

OpenStudy (anonymous):

oh okay rhx

OpenStudy (loser66):

yw